OpenStudy (anonymous):

The surface area of this pyramid is one hundred eighty-nine square centimeters. The area of the square base is forty-nine square centimeters. What is the area of one triangular face of the lateral surface? Show the steps of your solution and explain your work

OpenStudy (ashleyisakitty):

(189 - 49) /4 = 140/4 140/4 = ?

OpenStudy (anonymous):

so the answered will be 35

OpenStudy (ashleyisakitty):

correct :)

OpenStudy (anonymous):

thanks